Manganese(IV) oxide reacts with aluminum to form elemental manganese and aluminum oxide: 3MnO2+4Al→3Mn+2Al2O3 What mass of Al is required to completely react with 30.0 g MnO2?
Answer #1

3MnO2+4Al→3Mn+2Al2O3

m = 30 g of MnO2

MW of MnO2 = 86.9368

mol = 30/86.9368 = 0.345

3 mol of MnO2 : 4 mol of Al

therefore we need

4/3*0.345 = 0.46 mol of Al

MW of Al = 26.9

mass = mol*MW = 0.46*26.9 = 12.37 grams of Al